# Predictive Models in Sequential Recommendations: Bridging Performance Laws with Data Quality Insights

### v4

@@ Abstract
Sequential Recommendation (SR) plays a critical role in predicting users’ sequential preferences. Despite its growing prominence in various industries, the increasing scale of SR models incurs substantial computational costs and unpredictability, challenging developers to manage resources efficiently. Under this predicament, Scaling Laws have achieved significant success by examining the loss as models scale up. However, there remains a disparity between loss and model performance, which is of greater concern in practical applications. Moreover, as data continues to expand, it incorporates repetitive and inefficient data. In response, we introduce the Performance Law for SR models, which aims to theoretically investigate and model the relationship between model performance and data quality. Specifically, we first fit the HR and NDCG metrics to transformer-based SR models. Subsequently, we propose Approximate Entropy (ApEn) to assess data quality, presenting a more nuanced approach compared to traditional data quantity metrics. Our method enables accurate predictions across various dataset scales and model sizes, demonstrating a strong correlation in large SR models and offering insights into achieving optimal performance for any given model configuration.
@@ 3.2.2 ApEn Extension on Recommendation
In our subsequent calculation of ApEn, we set the tolerance $r=0$. This decision is due to the unique nature of recommended items, where products with similar IDs can potentially convey entirely different meanings.
@@ 3.3 Theorem of Performance Law
Before accurately predicting model performance, we need to consider factors such as model decay as the model size increases indefinitely, the relationship between model performance and loss, and how data scale correlates with token count and dataset Approximate Entropy (ApEn). For the first aspect, we introduce a $\log(.)$ decay term when the model layer $H$ and embedding dimension $d_{emb}$ is large, as proven in Theorem 3.5. Regarding the second aspect, we assume a linear relationship between model performance and loss, expressed as $\text{Performance}=1-kL$, Performance = HR@10, NDCG@10, For the final point, we provide a theoretical proof in Theorem 3.3, establishing the data scale $D$ is bounded by $D=\frac{\#Tokens}{ApEn}$. Building on the aforementioned theorems, we ultimately fit the model’s performance into the following equation:
@@
where $D^{\prime}=\frac{\#Tokens}{ApEn}$ represents the data parameter, $N$ is the number of model layers, and $d_{emb}$ is the item embedding dimension. When controlling model size, we only adjust the number of layers and embedding dimension, which is why our formula includes only these two parameters. However, our theoretical framework can still be applied for analysis when other model parameters are modified. Below is the detailed proof of our performance law theory:
@@
$\ \ \ \ -\sum_{\mathrm{x\in X}}\sum_{\mathrm{y\in X}}\pi(\mathrm{x})\mathrm{p_{xy}}log(\mathrm{p_{xy}}).$
@@
With the aforementioned lemma established, we proceed to conduct a detailed analysis of data volume within the framework of the original Scaling Law. Importantly, the minimum encoding length L(C) provides a nuanced perspective by integrating data compression rates (sayood2017introduction; Yin et al., 2024b), thereby offering insights into both the scale and quality of the data. This refined approach permits a more thorough assessment of how data affects model performance. Guided by these insights, we propose the following theorem:
@@
$|U|L(C)\geq(\Sigma_{i=1}^{n}|s_{i}|)/ApEn(S)$ (12)
@@
$L(C)\geq H_{2}(\mathcal{S})=\sum_{\forall s_{i}}p(s_{i})\log_{2}\frac{1}{p(s_{i})},$ (13)
where $H_{2}(\mathcal{S})$ represents the entropy of the sequence in base 2. Meanwhile, from Lemma 3.1, it follows that:

### v6

# Optimizing Sequential Recommendation Models with Scaling Laws and Approximate Entropy

@@ Abstract
Scaling Laws have emerged as a powerful framework for understanding how model performance evolves as they increase in size, providing valuable insights for optimizing computational resources. In the realm of Sequential Recommendation (SR), which is pivotal for predicting users’ sequential preferences, these laws offer a lens through which to address the challenges posed by the scalability of SR models. However, the presence of structural and collaborative issues in recommender systems prevents the direct application of the Scaling Law (SL) in these systems. In response, we introduce the Performance Law for SR models, which aims to theoretically investigate and model the relationship between model performance and data quality. Specifically, we first fit the HR and NDCG metrics to transformer-based SR models. Subsequently, we propose Approximate Entropy (ApEn) to assess data quality, presenting a more nuanced approach compared to traditional data quantity metrics. Our method enables accurate predictions across various dataset scales and model sizes, demonstrating a strong correlation in large SR models and offering insights into achieving optimal performance for any given model configuration.
@@ 3.2.2 ApEn Extension on Recommendation
In our subsequent calculation of ApEn, we set the tolerance $r=0$. This decision is due to the unique nature of recommended items, where products with similar IDs can potentially convey entirely different meanings. From Equations above, it can be observed that Approximate Entropy $ApEn$ has a positive correlation with the model’s duplication rate, whereas conventional entropy tends to have a negative correlation with data duplication rate. Therefore, although $ApEn$ is referred to as entropy, its trend is opposite to traditional entropy. To avoid confusion, we use $ApEn^{\prime}=1/ApEn$ as the final measure of Approximate Entropy.
@@ 3.3 Theorem of Performance Law
Before accurately predicting model performance, we need to consider factors such as model decay as the model size increases indefinitely, the relationship between model performance and loss, and how data scale correlates with token count and dataset Approximate Entropy (ApEn). For the first aspect, we introduce a $\log(.)$ decay term when the model layer $H$ and embedding dimension $d_{emb}$ is large, as proven in Theorem 3.5. Regarding the second aspect, we assume a linear relationship between model performance and loss, expressed as $\text{Performance}=1-kL$, Performance = HR@10, NDCG@10, For the final point, we provide a theoretical proof in Theorem 3.3, establishing the data scale $D$ is bounded by $D=\#Tokens\cdot ApEn^{\prime}$. Building on the aforementioned theorems, we ultimately fit the model’s performance into the following equation:
@@
where $D^{\prime}=\#Tokens\cdot ApEn^{\prime}$ represents the data parameter, $N$ is the number of model layers, and $d_{emb}$ is the item embedding dimension. When controlling model size, we only adjust the number of layers and embedding dimension, which is why our formula includes only these two parameters. However, our theoretical framework can still be applied for analysis when other model parameters are modified. Below is the detailed proof of our performance law theory:
@@
$\ \ \ \ =-\sum_{\mathrm{x\in X}}\sum_{\mathrm{y\in X}}\pi(\mathrm{x})\mathrm{p_{xy}}log(\mathrm{p_{xy}}).$
@@
With the aforementioned lemma established, we proceed to conduct a detailed analysis of data volume within the framework of the original Scaling Law. Importantly, the minimum encoding length L(C) provides a nuanced perspective by integrating data compression rates (Sayood, 2017; Yin et al., 2024b), thereby offering insights into both the scale and quality of the data. This refined approach permits a more thorough assessment of how data affects model performance. Guided by these insights, we propose the following theorem:
@@
$|U|L(C)\geq(\Sigma_{i=1}^{n}|s_{i}|)\cdot ApEn^{\prime}$ (12)
@@
$L(C)=\sum_{\forall s_{i}}p(s_{i})l_{i}=\sum_{\forall s_{i}}p(s_{i})log_{2}\frac{1}{2^{-l_{i}}}$ (13)
$\geq H_{2}(\mathcal{S})=\sum_{\forall s_{i}}p(s_{i})\log_{2}\frac{1}{p(s_{i})},$
where $H_{2}(\mathcal{S})$ represents the entropy of the sequence in base 2, and Kraft’s inequality (Kraft, 1949) $\sum_{i=1}^{n}2^{-l_{i}}\leq 1$ was utilized. Meanwhile, from Lemma 3.1, it follows that: